# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

### v2

@@ C.2.2 Simulation in Section 5
Table 2: The proportion of keep topic, ambiguous, and change of topic with varying $L=\{4,6,8,10,12,14,16,18\}$.
@@
C.2.3 Additional experiments
@@
C.3 Numerical Analysis of each Scenario in Figure 4
@@
Topic continuity. In Fig. 9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$. Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3, the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$. Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$. In the mixed-topics scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$. Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$. Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
@@
Ambiguous sequence. Input sequence $\bf X$ in Fig. 9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$. The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3. The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$. Then $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability. ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3. Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$. Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$. Therefore, the sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ is ambiguous, based on the Definition 4.
@@
Change of topic. For the input sequence $\bf X$ in Fig. 9(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 3: ${\bf e}_{5}>{\bf e}_{1}$. With ${\bf W}_{a}(\tau)$ trained in $\mathbb{T}_{a}$, the token ${\bf e}_{5}$ has $a_{1}=0.70$ and the token ${\bf e}_{4}$ has $a_{2}=a_{3}=a_{4}=0.10$. Obviously, $1\times 0.70=[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}>[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.10$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ consists of ${\bf e}_{5}$. However, $\widehat{\mathcal{G}}_{ab}^{(4)}$ consists of ${\bf e}_{4}$ instead of ${\bf e}_{5}$, due to $1\times 0.40=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{5}}<[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.20$. Since $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$ and $\widehat{\mathcal{G}}_{a}^{(4)}\not\subset\widehat{\mathcal{G}}_{ab}^{(4)}$, ${\bf W}_{ab}$ changes topic for the input sequence $\bf X$. Moreover, we have $({\bf e}_{5}\Rightarrow{\bf e}_{i})\in\mathcal{G}^{(4)}_{ab}$ for $i\in[4]$, as shown in Figure 3. Thus, the highest priority SCC (Definition 1) in $\mathbb{T}_{ab}$ is $\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. In the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, the lower-priority token ${\bf e}_{4}\not\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$ appears more frequently than the higher-priority token ${\bf e}_{5}\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$, illustrating our Theorem 4.

### v3

@@ D.3 Simulation in Section 5
Table 2: Proportion of keep topic, ambiguous, and change of topic with varying $L=\{4,6,8,10,12,14,16,18\}$.
@@
D.4 Additional Experiments for Convergence in Mixed-topics
@@
D.5 Numerical Analysis for Each Scenario in Figure 3
@@
Topic continuity. In Fig. 9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$. Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left), the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$. Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$. In the mixed-topics scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$. Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$. Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
@@
Ambiguous sequence. Input sequence $\bf X$ in Fig. 9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$. The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left). The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$. Then $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability. ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3 (left). Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$. Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$. Therefore, the sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ is ambiguous, based on the Definition 4.
@@
Change of topic. For the input sequence $\bf X$ in Fig. 9(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 3 (left): ${\bf e}_{5}>{\bf e}_{1}$. With ${\bf W}_{a}(\tau)$ trained in $\mathbb{T}_{a}$, the token ${\bf e}_{5}$ has $a_{1}=0.70$ and the token ${\bf e}_{4}$ has $a_{2}=a_{3}=a_{4}=0.10$. Obviously, $1\times 0.70=[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}>[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.10$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ consists of ${\bf e}_{5}$. However, $\widehat{\mathcal{G}}_{ab}^{(4)}$ consists of ${\bf e}_{4}$ instead of ${\bf e}_{5}$, due to $1\times 0.40=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{5}}<[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.20$. Since $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$ and $\widehat{\mathcal{G}}_{a}^{(4)}\not\subset\widehat{\mathcal{G}}_{ab}^{(4)}$, ${\bf W}_{ab}$ changes topic for the input sequence $\bf X$. Moreover, we have $({\bf e}_{5}\Rightarrow{\bf e}_{i})\in\mathcal{G}^{(4)}_{ab}$ for $i\in[4]$, as shown in Figure 3 (left). Thus, the highest priority SCC (Definition 1) in $\mathbb{T}_{ab}$ is $\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. In the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, the lower-priority token ${\bf e}_{4}\not\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$ appears more frequently than the higher-priority token ${\bf e}_{5}\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$, illustrating our Theorem 4.
Appendix E Experimental Details in Section 6
In this section, we provide the experimental details in four LLMs: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3. Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract as follows:
  1. 1.
    Extract the first $T$ words from paper $A$’s abstract as the input segment $\bf{X}$ from Topic $A$.
  2. 2.
    Randomly select 5 different papers from paper $A$, as papers in $\{B_{i}\}_{i=1}^{5}$.
  3. 3.
    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper $A$ as the the knowledge $A$, denoted as $\text{Ref}_{A}$. Each chunk is with 800 tokens.
  4. 4.
    Similarly, retrieve top 3 relevant excerpts from paper $B_{i}$ as the knowledge $B_{i}$, denoted as $\text{Ref}_{B_{i}}$.
  5. 5.
    Combine the knowledge from Topic $A$ and from Topic $B_{i}$ as the knowledge $AB_{i}$ for mixed-Topic, denoted as $\text{Ref}_{AB_{i}}$.
  6. 6.
    For the input segment $\bf{X}$, each LLM with the follow prompts, $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively. Notably, the only difference between $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$ is the reference excerpts provided $\text{Ref}_{A}$ or $\text{Ref}_{AB_{i}}$. All LLMs were queried with a temperature of 0 to match the greedy decoding in our theoretical framework. The maximum completion length was set to 1000 tokens to ensure that the generated continuations could complete the abstract.
    1. (a)
      $\text{Prompt}_{A}$:
      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{A}$. Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$. Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
    2. (b)
      $\text{Prompt}_{AB_{i}}$:
      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{AB_{i}}$. Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$. Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
  7. 7.
    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper $A$ and paper $B_{i}$.
E.1 Impact of the Input Length
To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$, as shown on the x-axis from Figure 5.
E.2 Impact of the Topic Ambiguity
We quantify topic (paper) ambiguity by the average similarity among each paper’s keywords. As arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs. To investigate the topic ambiguity, we fix the input length with $T=80$ for every paper as Topic $A$ and order papers by the average keywords similarity, as shown on the x-axis of Figure 5. Higher keywords similarity corresponds to lower topic ambiguity.
Appendix F Computational Resources for Experiments
In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed-topic, followed by 2 additional hours for next-token prediction.
In our experiments on LLMs, we query GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3 through API calls. All experiments were conducted on a standard laptop without specialized hardware. For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 20 hours of runtime, with a total of 13 million input tokens and 2.5 million output tokens. The total API usage cost for the experiments is approximately 200 USD.
Appendix G Impact Statement
Our investigation highlights fundamental differences between spontaneous topic changes in LLMs and spontaneous human thought, informing the development of more natural and flexible AI systems in domains such as customer service and mental health support. However, improving such capabilities can raise ethical considerations, including inadvertent manipulation of user focus, especially in persuasive or sensitive contexts. Our work, while largely theoretical, emphasizes the importance of fairness, privacy, and user autonomy as developers refine these systems to serve users’ interests, respect contextual boundaries, and remain accountable. This research has the potential to advance both Machine Learning and Human-Computer Interaction by informing new architectures that mimic human-like topic shifts; nevertheless, any real-world application of these findings should be accompanied by vigilant oversight to mitigate risks of misuse—such as deceptive or manipulative dialogue shifting. There are many other potential societal consequences of our work, none which we feel must be specifically highlighted here.